# Infrared Small Target Detection in Satellite Videos: A New Dataset and A Novel Recurrent Feature Refinement Framework

### v2

@@ Abstract
Multi-frame infrared small target (MIRST) detection in satellite videos is a long-standing, fundamental yet challenging task for decades, and the challenges can be summarized as: First, extremely small target size, highly complex clutters & noises, various satellite motions result in limited feature representation, high false alarms, and difficult motion analyses. Second, the lack of large-scale public available MIRST dataset in satellite videos greatly hinders the algorithm development. To address the aforementioned challenges, in this paper, we first build a large-scale dataset for MIRST detection in satellite videos (namely IRSatVideo-LEO), and then develop a recurrent feature refinement (RFR) framework as the baseline method. Specifically, IRSatVideo-LEO is a semi-simulated dataset with synthesized satellite motion, target appearance, trajectory and intensity, which can provide a standard toolbox for satellite video generation and a reliable evaluation platform to facilitate the algorithm development. For baseline method, RFR is proposed to be equipped with existing powerful CNN-based methods for long-term temporal dependency exploitation and integrated motion compensation & MIRST detection. Specifically, a pyramid deformable alignment (PDA) module and a temporal-spatial-frequency modulation (TSFM) module are proposed to achieve effective and efficient feature alignment, propagation, aggregation and refinement. Extensive experiments have been conducted to demonstrate the effectiveness and superiority of our scheme. The comparative results show that ResUNet equipped with RFR outperforms the state-of-the-art MIRST detection methods. Dataset and code are released at https://github.com/XinyiYing/RFR.
@@ I Introduction
Multi-frame infrared small target (MIRST) detection in satellite videos is a long-standing, fundamental yet challenging task in satellite surveillance systems, and has consecutively received considerable attention. Specifically, MIRST detection aims at localizing a scarce of candidate target pixels from image sequences captured by diverse earth-orbiting satellites, such as low earth-orbiting (LEO) satellites of 400-2K km, middle earth-orbiting (MEO) satellites of 2K-36K km, and geostationary earth-orbiting (GEO) satellites of 36K km. Compared to IRST detection in aerial-based and land-based imaging systems (mostly no more than 10 km), IRST detection in satellite videos remains more challenges, and has several unique characteristics.
@@
  • Various compound satellite motion: Translation, pitch, yaw, roll, jitter and scheduling of satellite largely reduce the imaging quality, resulting in weaker target intensity and blurred target contour. In addition, the coupled target and background motion largely increase the difficulty of motion extraction and exploitation.
@@
To address the aforementioned challenges, we build the first large-scale MIRST dataset in satellite videos (namely IRSatVideo-LEO dataset), including 200 sequences, 91366 frames with mask annotations. Due to limited availability of infrared satellite videos, IRSatVideo-LEO is a semi-simulated dataset with a real satellite image and synthesized satellite motion, target appearance, trajectory and intensity, which can provide a standard and efficient toolbox for satellite videos generation, and offer a reliable evaluation platform to facilitate the algorithms development. Note that, in this paper, we mainly focus on LEO and GEO satellites with smoothly moving satellite platforms, while the GEO satellite videos can also be simulated by our toolbox.
Based on the IRSatVideo-LEO dataset, we develop a recurrent feature refinement (RFR) framework as the baseline method for MIRST in satellite videos, which can be equipped with existing deep learning-based single-frame infrared small target (SIRST) detection methods for joint motion compensation and MIRST detection in a data-driven manner. Specifically, RFR employs recurrent framework to fully exploit the long-term temporal dependency from the entire input video. Through RFR, features are iteratively propagated and aggregated, which retains context information from all history memory to refine the current state. For feature propagation with moving platform, we employ deformable convolutions [8, 9, 10] to perform implicit motion compensation. Note that, considering the scale differences between background motion and target motion, we use a pyramid structure to decompose the motion, and perform target and background motion compensation in a coarse-to-fine manner. For feature aggregation, we propose a temporal-spatial-frequency modulation (TSFM) module to aggregate the beneficial temporal information from aligned features, and adaptively enhance the spatial and frequency information especially for small targets. In summary, the contributions of this paper can be summarized as follows.
  • We develop the first large-scale MIRST dataset in satellite videos, namely IRSatVideo-LEO to advance the technique development of MIRST detection in satellite videos.
@@ II-A Algorithms for IRST Detection
Based on the aforementioned single-frame methods, temporal clues within sequence images can be exploited [34, 20, 35, 5] to address blur, distortion and low-contrast targets within a single frame, and can offer reasonable history tracking and future forecasting. Specifically, a variety of spatio-temporal filters (e.g., 3-D matched filter [34, 36, 37], correlation filter [38, 39], particle filter [40, 41], wavelet filter [42, 43, 44]) are proposed to suppress low-frequency background and extract the region of interest. Recently, incorporating spatio-temporal information into local contrast-based methods [45, 46, 47] and low rank-based methods [22, 20, 48, 5, 49, 50] have raised more and more attention. Although these series of methods perform robust to low-contrast targets, their handcrafted features, fixed hyper-parameters and high computational complexity cannot generalize well to real scenes with complex background clutter and real-time application requirements. To address this problem, powerful CNN-based methods [35, 51, 52, 53] are emerged to learn trainable features in a data-driven manner, which can achieve state-of-the-art performance with high efficiency. Specifically, Li et. al. designed a direction-coded convolution block to distinguish the motion between targets and clutters. Zhang et. al. incorporated a spatial-temporal tensor (STT) optimization model into CNN-based method to achieve IRSTD in a model & data-driven manner. Yan et. al. proposed a spatio-temporal differential multi-scale attention network (STDMANet) to extract temporal multi-scale feature for IRSTD. However, these methods only focus on target motion under a relatively static field of view (some are with sudden movement by turnable collection), which cannot generalize well to complex background motion of satellite videos.
@@
Figure 1: Implementation details of the IRSatVideo-LEO dataset that consists of four steps: (a) Background sequence simulation. 3D attitude sequence is first used to generate global background sequence from a SWIR satellite image by homography transformation, and 2D location is then used to crop the local background sequence. (b) Target appearance sequence simulation. Several key target appearances (e.g., red and blue Gaussian kernel images) are used to generate target appearance sequence by quadratic interpolation1 (e.g., light red and light blue Gaussian kernel images are interpolated results). (c) Target trajectory simulation. We first generate the trajectory in the $1^{st}$ frame reference by connecting and smoothing multiple separate trajectories, and then employ satellite motion to perform reference transformation to generate trajectory in the current frame reference. (d) Target background superposition. Target appearance sequence and intensity are multiplied, which is then adaptive weighted summed by background sequence using target trajectory to generate the simulated sequence.
@@ II-B Datasets for IRST Detection
As the pioneering open-source SIRST detection datasets, Wang et. al. [23] synthesized 10100 images with real scenes and simulated targets. Then, Dai et. al. [24] collected and annotated 427 images to release the first real dataset (namely SIRST), which was then extended to SIRST-v2 with 514 images [61] . To further enlarge the available dataset, simulated NUDT-SIRST dataset and real IRSTD-1K dataset are proposed for over a thousand images. Besides the aforementioned land-based and aerial-based SIRST datasets, Wu et. al. [56] proposed the first space-based near-infrared tiny ship dataset mounted by low Earth-orbiting satellite.
Compared with SIRST detection datasets only provide appearance information, consecutive frames of MIRST detection datasets allow for a more comprehensive understanding of target behavior by additional motion patterns. In addition, the inherent appearance & radiation variations, motion blur, dynamic background and various annotation forms (e.g., point, box and mask) of MIRST detection datasets significantly increase the difficulty of this task. Hui et. al. [57] and Fu et. al. [58] released two real MIRST detection datasets through the 2nd and 3rd Sky Cup competitions, which contain 22 and 87 sequences with point and box annotations, respectively. Sun et. al. [59] semi-simulated 350 image sequences (namely SIATD) with background sequences captured by UAVs and simulated target appearance and motion. SIATD only provide point annotations. Sun et. al. [60] built a large-scale infrared small dim target dataset (IRDST). IRDST contains a real subset of 85 sequences and a simulated subset of 316 sequences with pixel-level annotations. Li et. al. [35] synthesized a multi-frame infrared small and dim target dataset (NUDT-MIRSDT) to explore MIRST detection under extremely low signal-to-noise ratio (i.e., SNR$<$3).
Despite the valuable contributions of open-source datasets in the field of MIRST detection, these datasets are all captured by aerial-based and land-based imaging systems, and have large differences against space-based datasets (see section I for details). Due to the scarcity and unavailability of infrared satellite videos, we decide to develop the first large-scale MIRST dataset in satellite videos, namely IRSatVideo-LEO dataset to advance the technique development of space-based MIRST detection.
@@ III The IRSatVideo-LEO Dataset
Large quantity, rich diversity, and high-quality labeled data is the cornerstone of deep learning-based algorithms, especially with the emergence of large-scale models. However, the inherent data scarcity of infrared satellite videos and high confidentiality of targets (e.g., aircraft, missile, and launch vehicle) greatly hinder the development of MIRST detection on satellite videos. Inspired by other data-scarcity fields [59, 62, 63, 64, 65, 66, 67], we develop a large-scale dataset (i.e., IRSatVideo-LEO) using semi-simulated implementation with a real satellite image and synthesized satellite motion, target appearance, trajectory and intensity. Please refer to section III-B for more details.
@@ III-A Background Image Collection
We collect 200 background images captured by the 7th band (i.e., SWIR 2.1-2.3 $\upmu m$) of Landset 8 and 9. To ensure the generalization of the dataset, we randomly sample locations across each continent and ocean on Earth, and the cloud cover ratios range from 0 to 61.25%. Figure 2 shows the cloud cover ratios and location distributions of IRSatVideo-LEO dataset.
@@ III-B1 Background Sequence Simulation
Figure 3: Illustrations of infrared small target and local background neighborhood. (a) Example images of infrared small targets. (b1) Local background neighborhood is extended from the BBox of target ($h_{0}$ in height and $w_{0}$ width) by $d$ in both height and width. (b2) illustrates the target region (i.e., yellow area) and local background region (i.e., purple area).
For 3D satellite locations, translation along the z-axis is ignored due to space-based imaging system. To generate the 2D satellite location sequence $(x_{t}^{B},y_{t}^{B})$, we randomly set the initial $(x_{1}^{B},y_{1}^{B})$ and final $(x_{T}^{B},y_{T}^{B})$ 2D locations, and employ quadratic interpolation [^1] Given the 2D initial $(x_{1},y_{1})$ and final $(x_{T},y_{T})$ points, we randomly extract a segment of a low-order curve with order $p\in(0,3)$, length $x_{T}-x_{1}$ along x-axis, and $y_{T}-y_{1}$ along y-axis. for intermediate locations. Note that, the velocities along x-axis and y-axis $v_{x}^{B}$=$\frac{x_{T}^{B}-x_{1}^{B}}{T}$, $v_{y}^{B}$=$\frac{y_{T}^{B}-y_{1}^{B}}{T}$ are randomly sampled from $[1/20,2]$. Then local background sequence is cropped according to the 2D satellite location sequence, and the pre-defined field of view. The process can be formulated as:
@@
Figure 4: Example target trajectories in satellite videos. (a), (b), (c) shows the trajectory of target without (w/o) and with (w.) one and two swerving. Note that, discrete points of target trajectory with a sampling rate of 20 are shown for better visualization. Density of discrete points represents the velocity of the target, and dense points represent high velocity. Red arrows are specific to the swerving position.
@@
$x_{t}^{B}$ Translation along x axis at time t Quadratic $v_{x}^{B}=(x_{T}^{B}-x_{1}^{B})/T$, $v_{x}^{B}\in$$[1/20,2]$
$y_{t}^{B}$ Translation along y axis at time t Quadratic $v_{y}^{B}=(y_{T}^{B}-y_{1}^{B})/T$, $v_{y}^{B}\in$$[1/20,2]$
$z_{t}^{B}$ Translation along z axis at time t Constant 00
@@
TABLE II: Statical comparisons among existing SIRST and MIRST detection datasets and our IRSatVideo-LEO dataset. “Seq.” and “Frame” represent the number of sequences and frames. “T-Num.”, “T-Size”, “T-SCR” represent the target number, average target size and average target SCR, respectively. Note that, “T-Size” is the pixel number with mask annotation, while is the box area with BBox annotation. “T-SCR” with BBox or point annotation () is calculated by the average value of SCR and SNR (i.e., replacing $\mu_{t}$ in Eq. 8 by the maximum value in target region).

### v3

@@ Abstract
Multi-frame infrared small target (MIRST) detection in satellite videos has been a long-standing, fundamental yet challenging task for decades, and the challenges can be summarized as follows: First, extremely small target size, highly complex clutters & noises, various satellite motions result in limited feature representation, high false alarms, and difficult motion analyses. In addition, existing methods are primarily designed for static or slightly adjusted perspectives captured by short-distance platforms, which cannot generalize well to complex background motion in satellite videos. Second, the lack of large-scale publicly available MIRST dataset in satellite videos greatly hinders the algorithm development. To address the aforementioned challenges, in this paper, we first build a large-scale dataset for MIRST detection in satellite videos (namely IRSatVideo-LEO), and then develop a recurrent feature refinement (RFR) framework as the baseline method for satellite motion estimation and compensation. Specifically, IRSatVideo-LEO is a semi-simulated dataset with synthesized satellite motion, target appearance, trajectory and intensity, which can provide a standard toolbox for satellite video generation and a reliable evaluation platform to facilitate the algorithm development. For the baseline method, RFR is proposed to be equipped with existing powerful CNN-based methods for long-term temporal dependency exploitation and integrated motion compensation & MIRST detection. Specifically, a pyramid deformable alignment (PDA) module is proposed to achieve effective feature alignment, and a temporal-spatial-frequency modulation (TSFM) module is proposed to achieve efficient feature aggregation and enhancement. Extensive experiments have been conducted to demonstrate the effectiveness and superiority of our scheme. The comparative results show that ResUNet equipped with RFR outperforms the state-of-the-art MIRST detection methods. Dataset and code are released at https://github.com/XinyiYing/RFR.
@@ I Introduction
Multi-frame infrared small target detection in satellite videos (MIRST-SatVideo) is a long-standing, fundamental yet challenging task in satellite surveillance systems, and has consecutively received considerable attention. Specifically, MIRST detection aims at localizing a scarce of candidate target pixels from image sequences captured by diverse earth-orbiting satellites, such as low earth-orbiting (LEO) satellites of 400-2K km, middle earth-orbiting (MEO) satellites of 2K-36K km, and geostationary earth-orbiting (GEO) satellites of 36K km. Compared to IRST detection in aerial-based and land-based imaging systems (mostly no more than 10 km), IRST detection in satellite videos remains more challenging, and has several unique characteristics.
@@
  • Various compound satellite motion: The presence of translation, pitch, yaw, roll, jitter, and satellite scheduling significantly degrades imaging quality, leading to diminished target intensity and blurred target outlines. In addition, the coupled target and background motion pose challenges for motion information extraction and utilization. Since existing methods mainly focus on scenarios with a relatively static or slightly modified field of view, they encounter difficulties in handling the sophisticated motion characteristics of satellite video.
@@
Dataset is the foundation underlying research development. Therefore, we build the first large-scale MIRST dataset in satellite videos (namely the IRSatVideo-LEO dataset), including 200 sequences, 91366 frames with mask annotations. Due to the limited availability of infrared satellite videos, IRSatVideo-LEO is a semi-simulated dataset with a real satellite image and synthesized satellite motion, target appearance, trajectory and intensity, which can provide a standard and efficient toolbox for satellite video generation, and offers a reliable evaluation platform to facilitate the algorithms development. Note that, in this paper, we mainly focus on LEO and GEO satellites with smoothly moving satellite platforms, while the GEO satellite videos can also be simulated by our toolbox.
Based on the IRSatVideo-LEO dataset, we develop a recurrent feature refinement (RFR) framework as the baseline method for MIRST in satellite videos. RFR can be equipped with existing deep learning-based single-frame infrared small target (SIRST) detection methods for joint satellite motion compensation and MIRST detection in a data-driven manner. Specifically, RFR employs recurrent framework [8, 9] to fully exploit the long-term temporal dependency from the entire input video. Through RFR, features are iteratively propagated and aggregated, which retains context information from all history memory to refine the current state. For feature propagation with moving platform, we employ deformable convolutions [10, 11, 12] to perform implicit motion compensation. Note that, considering the scale differences between background motion and target motion, we use a pyramid structure to decompose the motion, and perform target and background motion compensation in a coarse-to-fine manner. For feature aggregation, we propose a temporal-spatial-frequency modulation (TSFM) module to aggregate the beneficial temporal information from aligned features, and adaptively enhance the spatial and frequency information especially for small targets. In summary, the contributions of this paper can be summarized as follows:
  • This paper, for the first time, discusses the multi-frame infrared small target detection in satellite videos (MIRST-SatVideo), and builds the first large-scale dataset (namely IRSatVideo-LEO) to lay the foundation of research development.
@@ II-A Algorithms for IRST Detection
Based on the aforementioned single-frame methods, temporal clues within sequence images can be exploited [36, 22, 37, 5] to address blur, distortion and low-contrast targets within a single frame, and can offer reasonable history tracking and future forecasting. Specifically, a variety of spatio-temporal filters (e.g., 3-D matched filter [36, 38, 39], correlation filter [40, 41], particle filter [42, 43], wavelet filter [44, 45, 46]) are proposed to suppress low-frequency background and extract the region of interest. Recently, incorporating spatio-temporal information into local contrast-based methods [47, 48, 49] and low rank-based methods [24, 22, 50, 5, 51, 52] have raised more and more attention. Although these series of methods perform robust to low-contrast targets, their handcrafted features, fixed hyper-parameters and high computational complexity cannot generalize well to real scenes with complex background clutter and real-time application requirements. To address this problem, powerful CNN-based methods [37, 53, 54, 55] have emerged to learn trainable features in a data-driven manner, which can achieve state-of-the-art performance with high efficiency. Specifically, Li et. al. designed a direction-coded convolution block to distinguish the motion between targets and clutters. Zhang et. al. incorporated a spatial-temporal tensor (STT) optimization model into CNN-based method to achieve IRSTD in a model & data-driven manner. Yan et. al. proposed a spatio-temporal differential multi-scale attention network (STDMANet) to extract temporal multi-scale features for IRSTD. However, these methods only focus on target motion under a relatively static field of view (some are with sudden movement by turnable collection), which cannot generalize well to the complex background motion of satellite videos.
Recently, motion estimation and compensation have been widely used in many video processing tasks, such as video super-resolution [56, 57, 58], video inpainting [59, 60], video object detection [61, 62, 2]. Compared with two-step optical flow-based approaches [57, 59, 61] (i.e., motion estimation by optical flow approaches and frame alignment by warping) that always leads to ambiguity and duplication, deformable convolution-based methods [58, 60] achieve motion compensation in a unified step by dynamically adjusted sampling grid in a learnable manner. Therefore, we employ deformable convolution for satellite motion compensation in our paper.
@@
Figure 1: Implementation details of the IRSatVideo-LEO dataset that consists of four steps: (a) Background sequence simulation. 3D attitude sequence is first used to generate a global background sequence from a SWIR satellite image by homography transformation, and 2D location is then used to crop the local background sequence. (b) Target appearance sequence simulation. Several key target appearances (e.g., red and blue Gaussian kernel images) are used to generate target appearance sequence by quadratic interpolation1 (e.g., light red and light blue Gaussian kernel images are interpolated results). (c) Target trajectory simulation. We first generate the trajectory in the $1^{st}$ frame reference by connecting and smoothing multiple separate trajectories, and then employ satellite motion to perform reference transformation to generate trajectory in the current frame reference. (d) Target background superposition. Target appearance sequence and intensity are multiplied, which is then adaptive weighted summed by background sequence using target trajectory to generate the simulated sequence.
@@ II-B Datasets for IRST Detection
As the pioneering open-source SIRST detection datasets, Wang et. al. [25] synthesized 10100 images with real scenes and simulated targets. Then, Dai et. al. [26] collected and annotated 427 images to release the first real dataset (namely SIRST), which was then extended to SIRST-v2 with 514 images [70]. To further enlarge the available dataset, the simulated NUDT-SIRST dataset and real IRSTD-1K dataset are proposed for over a thousand images. Besides the aforementioned land-based and aerial-based SIRST datasets, Wu et. al. [65] proposed the first space-based near-infrared tiny ship dataset mounted by a low Earth-orbiting satellite.
Compared with SIRST detection datasets that only provide appearance information, consecutive frames of MIRST detection datasets allow for a more comprehensive understanding of target behavior by additional motion patterns. In addition, the inherent appearance & radiation variations, motion blur, dynamic background and various annotation forms (e.g., point, box and mask) of MIRST detection datasets significantly increase the difficulty of this task. Hui et. al. [66] and Fu et. al. [67] released two real MIRST detection datasets through the 2nd and 3rd Sky Cup competitions, which contain 22 and 87 sequences with point and box annotations, respectively. Sun et. al. [68] semi-simulated 350 image sequences (namely SIATD) with background sequences captured by UAVs and simulated target appearance and motion. SIATD only provides point annotations. Sun et. al. [69] built a large-scale infrared small dim target dataset (IRDST). IRDST contains a real subset of 85 sequences and a simulated subset of 316 sequences with pixel-level annotations. Li et. al. [37] synthesized a multi-frame infrared small and dim target dataset (NUDT-MIRSDT) to explore MIRST detection under extremely low signal-to-noise ratio (i.e., SNR$<$3).
Despite the valuable contributions of open-source datasets in the field of MIRST detection, these datasets are all captured by aerial-based and land-based imaging systems, and have large differences against space-based datasets (see section I for details). Due to the scarcity and unavailability of infrared satellite videos, we decide to develop the first large-scale MIRST dataset in satellite videos, namely the IRSatVideo-LEO dataset, to advance the technique development of space-based MIRST detection.
@@ III The IRSatVideo-LEO Dataset
Large quantity, rich diversity, and high-quality labeled data are the cornerstone of deep learning-based algorithms, especially with the emergence of large-scale models. However, the inherent data scarcity of infrared satellite videos and high confidentiality of targets (e.g., aircraft, missile, and launch vehicle) greatly hinder the development of MIRST detection on satellite videos. Inspired by other data-scarcity fields [68, 71, 72, 73, 74, 75], we develop a large-scale dataset (i.e., IRSatVideo-LEO) using semi-simulated implementation with a real satellite image and synthesized satellite motion, target appearance, trajectory and intensity. Please refer to section III-B for more details.
@@ III-A Background Image Collection
We collect 200 background images captured by the 7th band (i.e., SWIR 2.1-2.3 $\upmu m$) of Landsat 8 and 9. To ensure the generalization of the dataset, we randomly sample locations across each continent and ocean on Earth, and the cloud cover ratios range from 0 to 61.25%. Figure 2 shows the cloud cover ratios and the location distributions of the IRSatVideo-LEO dataset.
@@ III-B1 Background Sequence Simulation
Figure 3: Illustrations of infrared small target and local background neighborhood. (a) Example images of infrared small targets. (b1) Local background neighborhood is extended from the BBox of the target ($h_{0}$ in height and $w_{0}$ width) by $d$ in both height and width. (b2) illustrates the target region (i.e., yellow area) and local background region (i.e., purple area).
For 3D satellite locations, translation along the z-axis is ignored due to the space-based imaging system. To generate the 2D satellite location sequence $(x_{t}^{B},y_{t}^{B})$, we randomly set the initial $(x_{1}^{B},y_{1}^{B})$ and final $(x_{T}^{B},y_{T}^{B})$ 2D locations, and employ quadratic interpolation [^1] Given the 2D initial $(x_{1},y_{1})$ and final $(x_{T},y_{T})$ points, we randomly extract a segment of a low-order curve with order $p\in(0,3)$, length $x_{T}-x_{1}$ along x-axis, and $y_{T}-y_{1}$ along y-axis. for intermediate locations. Note that, the velocities along x-axis and y-axis $v_{x}^{B}$=$\frac{x_{T}^{B}-x_{1}^{B}}{T}$, $v_{y}^{B}$=$\frac{y_{T}^{B}-y_{1}^{B}}{T}$ are randomly sampled from $[1/20,2]$. Then local background sequence is cropped according to the 2D satellite location sequence, and the pre-defined field of view. The process can be formulated as:
@@
Figure 4: Example target trajectories in satellite videos. (a), (b), (c) shows the trajectory of target without (w/o) and with (w.) one and two swerving. Note that, discrete points of target trajectory with a sampling rate of 20 are shown for better visualization. The density of discrete points represents the velocity of the target, and dense points represent high velocity. Red arrows are specific to the swerving position.
@@
$x_{t}^{B}$ Translation along x-axis at time t Quadratic $v_{x}^{B}=(x_{T}^{B}-x_{1}^{B})/T$, $v_{x}^{B}\in$$[1/20,2]$
$y_{t}^{B}$ Translation along y-axis at time t Quadratic $v_{y}^{B}=(y_{T}^{B}-y_{1}^{B})/T$, $v_{y}^{B}\in$$[1/20,2]$
$z_{t}^{B}$ Translation along z-axis at time t Constant 00
@@
TABLE II: Statical comparisons among existing SIRST and MIRST detection datasets and our IRSatVideo-LEO dataset. “Seq.” and “Frame” represent the number of sequences and frames. “T-Num.”, “T-Size” and “T-SCR” represent the target number, average target size and average target SCR, respectively. Note that, “T-Size” is the pixel number with mask annotation, while is the box area with BBox annotation. “T-SCR” with BBox or point annotation () is calculated by the average value of SCR and SNR (i.e., replacing $\mu_{t}$ in Eq. 8 by the maximum value in target region).